OpenStudy (anonymous):

Determine the molarity: 275ml of 0.16M NaOH added to 525ml of 0.040 M NaOH.

OpenStudy (chmvijay):

M1V1+M2V2=M(V1+V2) M=?

OpenStudy (chmvijay):

did u get it or NOOO

OpenStudy (chmvijay):

0.16*275+0.040*525=M(275+525)

OpenStudy (chmvijay):

find M=?